1. Conduct a χ2test for the following hypotheses at a 90% level of confidence:
Ho: fofits fe    (observed frequencies are consistent with those expected)
H1: fo does not fit fe   (observed frequncies are inconsistent with those expected)
Doneness f0 fe     
Rare 20 28 χ2=
Medium 40 35
Well 60 57 χ2α =
Ho: is... accepted/rejected
Does fofit fe?
Answer #1

H0: Null Hypothesis: Observed frequencies are consistent with those expected

H1: Alternative Hypothesis: Observed frequencies are not consistent with those expected

Test statistic \chi ^{2} is got as follows:

Doneness fo fe (fo - fe)2/fe
Rare 20 28 2.2857
Medium 40 35 0.7143
Well 60 57 0.1579
Total = \chi ^{2} = 3.1579

Degrees of Freedom = 3 - 1 = 2

\alpha = 0.10

From Table, critical value of \chi ^{2} = 4.605

H0 is accepted

fo fits fe